OpenStudy (anonymous):

Find the zeros of the function ,giving the multiplicity *PLEASE HELP* : f(x)=-3(x+1/2)(x-3)^3 ????

OpenStudy (anonymous):

Do you know what the zero-product property is?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

Alright here it is. If (x-5)(x-3) = 0 then (x-5) = 0 or (x-3) = 0

OpenStudy (anonymous):

Do you think you could do that with the problem you have above?

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

What do you do with the number in front of the parenthesis?

OpenStudy (anonymous):

Take each product and set it = to 0

OpenStudy (anonymous):

You can set that = to zero but -3≠0 so that won't be a solution

OpenStudy (anonymous):

Now set the other two equal to zero.

OpenStudy (anonymous):

-1/2 and 3...

OpenStudy (anonymous):

Yeah those are your two solutions. But you have to state the multiplicities. Do you know what they are?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

okay so x+1/2=0, x=-1/2 has a multiplicity of one and x-3=0 , x=3 has a multiplicity of 3. That's because it was cubed.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So do you get it now Miranda?

OpenStudy (anonymous):

Miranda?

OpenStudy (anonymous):

Nevermind.

OpenStudy (anonymous):

1 more?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

F(x)=x^3-2x^2+x

OpenStudy (anonymous):

Find zeros?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Okay. What can we factor out right now?

OpenStudy (anonymous):

What does each term have at least one of

OpenStudy (anonymous):

So: x(x^2-2x+1)?

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

Now can you factor what's inside the parenthesis?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

F(x)=x^3-2x^2+x

OpenStudy (anonymous):

1 more?

OpenStudy (anonymous):

We're not done yet. we have yet to find the zeros.

OpenStudy (anonymous):

X=1

OpenStudy (anonymous):

No

OpenStudy (anonymous):

factor what's inside the parenthesis.

OpenStudy (anonymous):

x(x^2-2x+1)?

OpenStudy (anonymous):

hint: You'll have x(x-?)(x-?)

OpenStudy (anonymous):

I did : (x-1)(x-1), set it equal to 0, its x=1,x=1

OpenStudy (anonymous):

Oh okay. Well you didn't tell me you factored it. Yeah that would be right. So in the end you would have x(x-1)(x-1) which can be simplified to x(x-1)^2 You told me one zero, what's the other one and also tell me the multiplicities.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Okay so you have to make sure you set all the products = to 0 You forgot the x that's standing all alone. so x = 0 or x-1=0

OpenStudy (anonymous):

x=0 has a multiplicity of 1 and x=1 has a mutliplicity of 2. I think you should get the multiplicity thing now. we had (x-1)(x-1) or (x-1)^2 That's two of the same zero, which is the multiplicity.

OpenStudy (anonymous):

K,thanks

OpenStudy (anonymous):

Are you good now? You're a pretty smart cookie.